Brad is half the age of his father, who is 3 times the age of his niece Imelda. How old is Brad when the three's combined age in years is that of Brad's 88 year old grandmother?
Answer by stanbon(75887)

Brad is half the age of his father, who is 3 times the age of his niece Imelda. How old is Brad when the three's combined age in years is that of Brad's 88 year old grandmother?
-------
Equations:
b = (1/2)f
f = 3n
b + f + n = 88
-----
Substitute for b and n and solve for "f":
(1/2)f + f + (1/3)f = 88
------
3f + 6f + 2f = 6*88
11f = 6*88
f = 6*8 = 48 yrs (father's age)
----
b = (1/2)f = 24 yrs (brad's age)
----
n = (1/3)f = 16 yrs (niece's age)
======================================
Cheers,
Stan H.
=================